# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

### 24 Sep 2025 v2

@@ 8. Findings: Rethinking Manager Clone Agent Design in Future Workplace (RQ3)
Having surfaced both hopes and anxieties around Manager Clone Agents, participants turned toward imagining how such systems should be designed and deployed. Three interconnected design orientations emerged: keeping workers’ needs at the center (see subsection 8.1), positioning Manager Clone Agents as supportive extensions rather than replacements of managers (see subsection 8.2), and structuring agents along graduated levels of autonomy to allow for context-sensitive use and gradual trust-building (see subsection 8.3). These directions point to a future where Manager Clone Agents are not simply more efficient delegators, but carefully crafted collaborators whose roles are explicitly delimited to protect and nurture human relationships in teams and organizations.
8.1. Worker-Oriented: Prioritizing Workers’ Need
Workers repeatedly highlighted that the introduction of Manager Clone Agents does not only affect managers but also reshapes their own everyday realities. While managers may see such systems as convenient extensions of their authority, workers are the ones who must interact with them directly and continuously, who are managed by those clone agents, and who directly bear the consequences of the use of managerial clone agents. As M10 put it bluntly, most workers approach their jobs “just to earn a wage,” and if a digital agent suddenly appeared to “constantly command” them, making work “increasingly difficult” or adding “many annoying tasks,” they would “hate it deeply.” In such cases, rather than lightening workloads, the Manager Clone Agents multiplies expectations, leaving workers more burdened than before. Therefore, it’s highly important to take worker’s pains and needs into consideration when designing Manager Clone Agents.
@@ 8.1. Worker-Oriented: Prioritizing Workers’ Need
Concerns about excessive control and surveillance were repeatedly emphasized. M11 observed that workers may feel as if they are reporting to “two leaders,” one being an extremely rational and emotionally detached version of their manager who “keeps asking endless questions.” M9, reflecting on her own workplace, stressed that she did not want colleagues to feel as if they were “always on the spot” or “watched the way that some companies just have your camera on all day.” For her, technology that fosters this feeling amounts to micromanagement.
Workers, positioned more vulnerably within workplace hierarchies, are especially attentive to whether Manager Clone Agents reinforce fairness and respect in their daily labor. W1, pointed to structural inequities that might emerge: “Workers willing to rely heavily on agents could offload routine labor and free up time for other pursuits, while those unwilling or unable to adopt the technology would be forced to shoulder more tasks themselves.” Over time, she argued, “such asymmetries could create divisions between groups of workers and amplify perceptions of unfairness.”
At the same time, participants highlighted that they care most about whether such systems deliver tangible value for workers. W3, for example, welcomed the idea that “An agent could improve and polish documents before sending them to the manager”. More broadly, as mentioned in section 6 workers expressed interest in Manager Clone Agents streamlining administrative tasks relaying feedback more consistently, etc. In these cases, the technology was seen as a supportive tool that reduced invisible labor.
In addition, whether Manager Clone Agents are perceived as supportive or extractive depends greatly on how their role is framed. The Manager Clone Agent’s role can be perceived differently as a tool or an assistant of the managers or the manager itself. As M12 summarized workplace dynamics often feel like “us versus them,” whether the Manager Clone Agents shows it cares about workers will shape how they respond to it.
@@
These tensions further underscore the importance of clarity and consent. Workers repeatedly pointed out that the legitimacy of Manager Clone Agents depends on explicit agreements about where they may or may not be used. For example, W6, a junior programmer, argued that a one-on-one meeting is a crucial opportunity to “leave a good impression” and demonstrate initiative, and substituting an agent in this setting would feel “disrespectful.” W8 echoed this sentiment, insisting that in contexts “where emotional intelligence is important, I want the real person.” At the same time, he envisioned negotiating expectations directly with his manager: deciding together which interactions could be delegated to the agent and which must remain human-led.
@@
Yet not all contexts lent themselves to easy consensus. M11, a Vice President of the company, described how he considered onboarding sessions fully replaceable: “It’s repetitive and a waste of time.” By contrast, W1, who’s only been working for 1-2 years, stressed that orientation was precisely when human presence mattered most: “If I’m a new employee joining a company, there might be many things that need to be coordinated with me, I would prefer a real person to handle that coordination with me.” Such disagreements illustrate how consent is not only about formal negotiation but also about uneven power and priorities. Managers may be motivated to offload repetitive duties, while workers may see those same moments as critical to relationship-building and trust. Because not all workers feel equally empowered to refuse a Manager Clone Agent’s presence, participants advocated for organizational safeguards, such as granting workers the protected right to request “human-only” interactions or limiting the contexts in which Manager Clone Agents may operate.
8.2. Revisit “Enhance vs. Replace”: Manager Clone Agents as Interpersonal Bond
@@
The HCI community has long advocated for AI that enhances rather than replaces human capabilities (Capel and Brereton, 2023). Our findings extend this debate by showing how the line between enhancement and replacement is inseparable from the quality of manager–worker relationships. Participants judged the value of Manager Clone Agents less by efficiency alone than by whether these systems strengthened or weakened the bonds between humans. Importantly, we found sometimes managers and workers often imagined enhancement in different terms, reflecting their divergent stakes in organizational life.
@@
For managers, enhancement meant receiving support that extended their capacity without threatening authority or identity. M9 expressed her wish for the agent to help her “be a better manager,” for instance by reminding her when she was “over budget” or helping her track priorities, rather than taking over leadership itself. Similarly, M11 stressed that the agent should remain in a service role, “approving a travel request if a boss is unavailable,” but not as “always pointing fingers.” M8 echoed this distinction, noting she was comfortable when the system was “clarifying what’s already there,” but grew uneasy when it began “creating new responses,” since that crossed into authorship and identity.
@@
Workers, by contrast, emphasized enhancement as a way to maintain access and grow. W3 imagined using a Manager Clone Agent to manage emotions before approaching her boss: when she feared she might “lose control” in a heated moment, the agent could help frame her message more constructively. W8 described the agent as “an extension of the manager” that could provide information or quick decisions while still leaving escalation to the human manager. For workers, enhancement thus meant greater availability and emotional scaffolding, rather than authority protection.
@@
Yet the line between enhancement and replacement remained blurry. W4’s story illustrated this dilemma: “I asked, ‘Then how should I communicate this?’ The Agent replied, ‘Do you want me to pass it on or do you want to do it yourself?’ I said, ‘What do you think?’ The Agent said, ‘If you do it, it can help you practice your communication skills, but you might not get what you want. But if I do it for you, it might be more effective but you wouldn’t be involved in the process.’” While workers saw this as a trade-off between empowerment and convenience, managers like M3 worried that such mediation of clone agents could “build an imaginary wall that ends up hurting both sides,” deepening misunderstanding rather than reducing it.
@@
Even as participants offered these visions of enhancement, they consistently stressed the importance of retaining human autonomy. M8 argued that preserving creativity and trust within teams is critical, noting that relationships, respect, and psychological safety “take time to develop” and cannot be delegated to machines. She warned that if technology undermines “any version of I can’t trust it, or I don’t know, or I don’t feel comfortable,” it risks eroding the interpersonal foundations that drive productivity. M9 echoed this sentiment by emphasizing the value of choice. She wanted to “see it all, but be able to say like, I’m not busy right now, but I trust that bot to do the answering for me or I’ll get back to you in 5 minutes.”
@@ 8.3. Tiered Autonomy: Gradual Trust-Building and Flexible Configuration
As participants imagined how Manager Clone Agents might operate in practice, they often turned to the metaphor of “levels of autonomy” familiar from self-driving cars (M2, M10). Rather than treating Manager Clone Agents as a single, fixed entity, they described autonomy as a spectrum, structured along dimensions such as representation, proactivity, and delegation. These dimensions offered a shared language for debating what forms of cloning would feel acceptable, what boundaries should be respected, and how much authority agents ought to exercise in different workplace contexts.
@@
Participants emphasized the spectrum of representational fidelity, ranging from highly stylized, cartoon-like avatars to hyper-realistic deepfakes indistinguishable from the human manager. For modality, At the lowest level, Manager Clone Agents may exist as text-based entities; at higher levels, they may incorporate voice, video, or even embodied presence in virtual or physical environments. While M13 expressed discomfort with Manager Clone Agents that “perfectly replicated their speech patterns and idiosyncrasies”, finding the imitation “uncanny and inauthentic”, workers similarly noted heightened anxiety when they could not discern whether communication originated from the manager or the agent (W2, W7). M3, W7 and W8 agreed that abstract forms of representation, such as “cartoon styled (M3) or “orb-like” voice interface (W7, W8), were perceived as less deceptive and more broadly acceptable. W8 further explained that embodiment affects comfort: “If it’s an avatar, it requires more attention, more effort to be present, like I’m performing extra socially.” Text-based agents, by contrast, were viewed as lightweight and flexible, more easily integrated into everyday practices such as Slack channels.
@@
Another dimension concerns the level of initiative that Manager Clone Agents should exercise, which participants saw as varying by task. At the lowest level, agents might simply observe or log interactions without intervening. At intermediate levels, they could relay pre-authorized statements or provide clarifications. The highest levels would involve spontaneous contributions or even steering discussions that many participants found unacceptable for sensitive or high-stakes contexts. As noted in subsection 8.2, M8 and M9 worried about being misrepresented by agents making statements they had not endorsed, or by discussions diverted in their absence in ways that later shaped decisions. M10 cautioned that disclosure must be carefully limited, warning, “The authority to disclose information should actually be tightened as well. For example, if it involves company finances and someone says that cash flow is tight, workers might panic.” Similarly, M11 argued that strong presence could backfire sometimes, suggesting, “Don’t make their presence too strong. Even just occasional reminders or emotional support are enough to help everyone gradually accept this new colleague. If their presence is too strong, everyone will feel threatened.”
@@
Finally, the scope of action delegated to Manager Clone Agents emerged as a central concern. Low-delegation configurations would confine agents to communication support, while high-delegation systems could independently execute managerial tasks, such as reallocating resources or authorizing workflows, without human intervention. M7 stressed limits: “As a strictly constrained ‘assistant,’ emphasize ‘reminders’ rather than ‘decisions,’ with transparency and explainability prioritized.” M10 described the risks of overstepping: “Some overly specific suggestions should be constrained. The remaining decision space should be left to the person doing the work. Otherwise, it’s like having a CTO constantly coaching right next to the coder, which makes workers uncomfortable.” M5 proposed explicit labeling: “Label the level of an authorized agent, for example, before getting a Gold level no one pays attention, but after getting Gold any command can be executed. So you can design agent tiers to let the interlocutor understand their current permissions.”

### 29 Jan 2026 v3

@@ 7. Findings (RQ3): Designing for Manager Clone Agents in Future Workplace
Having surfaced both hopes and anxieties around Manager Clone Agents, participants turned toward imagining how such systems should be designed and deployed. Three interconnected design orientations emerged: keeping workers’ needs at the center (see subsection 7.1), positioning Manager Clone Agents as supportive extensions rather than replacements for managers (see subsection 7.2), and structuring agents along graduated levels of autonomy to allow context-sensitive use and gradual trust-building (see subsection 7.3). Alongside these orientations, participants also ventured into more disruptive speculative imaginations that unsettled the assumed permanence of managerial roles, pointing toward alternative futures in which workers resist, managers fade, or clone agents are repurposed in unexpected ways (see subsection 7.4).
7.1. Worker-Oriented: Supporting Workers Rather Than Surveilling Them
While managers may see such systems as convenient extensions of their authority, workers are the ones who interact with them and who directly bear the consequences of managerial clone agents usage. The participants, therefore, pointed out that it is highly important to prioritize the pains and needs of the workers and support their work when designing Manager Clone Agents.
@@ 7.1. Worker-Oriented: Supporting Workers Rather Than Surveilling Them
Concerns about excessive control and surveillance were repeatedly emphasized. M19 observed that workers can feel as if they are reporting to “two leaders,” one being an extremely rational and emotionally detached version of their manager who “keeps asking endless questions.” M17, reflecting on her own workplace, stressed that she did not want colleagues to feel as if they were “always on the spot” or “watched the way that some companies just have your camera on all day.” For her, technology that fosters this feeling amounts to micromanagement.
Both managers and workers emphasized that their top concern is whether these systems provide concrete benefits to workers; otherwise, they will face pushback. W7, for example, proposed the idea that “An agent could improve and polish documents before sending them to the manager”. More broadly, as mentioned in section 5 workers expressed interest in Manager Clone Agents streamlining administrative tasks in team collaboration. In these cases, clone agents were seen as a supportive tool that reduced invisible labor. As M18 put it bluntly, most workers approach their jobs “just to earn a wage,” and if a digital agent suddenly appeared to “constantly command” them, making work “increasingly difficult” or adding “many annoying tasks,” they would “hate it deeply.” As M21 summarized: “workplace dynamics often feel like us versus them, whether the Manager Clone Agents shows that it cares about workers will shape how they respond to it.
@@
These tensions further underscore the importance of clarity and consent. Workers pointed out that the legitimacy of Manager Clone Agents depends on explicit agreements about where they may or may not be used. For example, W12, a junior programmer, argued that a one-on-one meeting is a crucial opportunity to “leave a good impression” and demonstrate initiative, and substituting an agent in this setting would feel “disrespectful.” W15 echoed this sentiment, insisting that in contexts “where emotional intelligence is important, I want the real person.” At the same time, he envisioned negotiating expectations directly with his manager: deciding together which interactions could be delegated to the agent and which must remain human-led.
@@
Yet not all contexts lend themselves to easy consensus. M19, a company Vice President, described how he considered on-boarding sessions fully replaceable: “It’s repetitive and a waste of time.” By contrast, W1, who has only worked for 1-2 years, stressed that orientation was precisely when human presence mattered most: “If I’m a new employee joining a company, there might be many things that need to be coordinated with me, I would prefer a real person to handle that coordination with me.” Such disagreements illustrate how consent is not only about formal negotiation, but also about uneven power and priorities. Managers may be motivated to offload repetitive tasks, while workers may see those same moments as critical to building relationships and trust. Because not all workers feel adequately empowered to refuse the presence of a Manager Clone Agent, the participants advocated organizational safeguards, such as granting workers the protected right to request “human-only” interactions or limiting the contexts in which Manager Clone Agents may operate.
7.2. Revisit “Enhance vs. Replace”: Bridging the Relationships Rather Than Breaking Them
@@
The HCI community has long advocated for AI that enhances rather than replaces human capabilities (Capel and Brereton, 2023). Our findings extend this debate by showing how the line between enhancement and replacement is inseparable from the quality of manager–worker relationships. Participants judged the value of Manager Clone Agents less by efficiency alone than by whether these systems strengthened or weakened the bonds between humans. Importantly, we found that sometimes managers and workers often imagined enhancement in different terms, reflecting their divergent values and stakes in organizational life.
@@
For managers, enhancement meant extending their capacity without threatening authority or identity. M17 expressed her wish for the agent to help her “be a better manager,” for example, by reminding her when she was “over budget” or helping her track priorities, rather than taking over leadership itself. Similarly, M19 stressed that the agent should remain in a service role, “approving travel requests when a boss is unavailable” but not “always pointing fingers.” M16 echoed this distinction, noting that she was comfortable when the system was “clarifying what’s already there,” but grew uneasy when it began “creating new responses,” since that crossed into authorship and identity.
@@
For workers, enhancement meant maintaining access to the leadership while buffering the potential conflicts. W7 imagined using a Manager Clone Agent to manage emotions before approaching her boss: when she feared she might “lose control” in a heated moment, the agent could help frame her message more constructively. W15 described the agent as “an extension of the manager” that could provide information or quick decisions while still leaving the escalation to the human manager. For workers, enhancement thus meant greater availability and emotional scaffolding, rather than authority protection.
@@
Yet the line between enhancement and replacement remained blurry. W8’s story illustrated this dilemma: “I asked, ‘Then how should I communicate this?’ The Agent replied, ‘Do you want me to pass it on or do you want to do it yourself?’ I said, ‘What do you think?’ The Agent said, ‘If you do it, it can help you practice your communication skills, but you might not get what you want. But if I do it for you, it might be more effective but you wouldn’t be involved in the process.’” While workers saw this as a trade-off between empowerment and convenience, managers like M5 worried that such mediation of clone agents could “build an imaginary wall that ends up hurting both sides,” deepening misunderstanding rather than reducing it.
@@
Even as participants offered these visions of enhancement, they consistently stressed the importance of retaining human autonomy. M16 argued that preserving creativity and trust within teams is critical, noting that relationships, respect, and psychological safety “take time to develop” and cannot be delegated to machines. She warned that if technology undermines “any version of I can’t trust it, or I don’t know, or I don’t feel comfortable,” it risks eroding the interpersonal foundations that drive productivity. M17 echoed this sentiment by emphasizing the value of choice. She wanted to “see it all, but be able to say like, I’m not busy right now, but I trust that bot to answer for me or I’ll get back to you in 5 minutes.”
@@ 7.3. Tiered Autonomy: Gradual Trust-Building and Flexible Configuration
As participants imagined how Manager Clone Agents might operate in practice, they often turned to the metaphor of “levels of autonomy” familiar from self-driving cars (M4, M18). Rather than treating Manager Clone Agents as a single, fixed entity, they described autonomy as a spectrum, structured along dimensions such as representation, proactivity, and delegation. These dimensions offered a shared language for debating what forms of cloning would feel acceptable, what boundaries should be respected, and how much authority agents ought to exercise in different workplace contexts.
@@
Participants emphasized the spectrum of representational fidelity, ranging from highly stylized, cartoon-like avatars to hyper-realistic deepfakes indistinguishable from the human manager. For modality, at the lowest level, Manager Clone Agents may exist as text-based entities; at higher levels, they may incorporate voice, video, or even embodied presence in virtual or physical environments. M22 expressed discomfort with Manager Clone Agents that “perfectly replicated their speech patterns and idiosyncrasies”, finding the imitation “uncanny and inauthentic”; workers similarly noted heightened anxiety when they could not discern whether communication originated from the manager or the agent (W3, W14). M5, W14 and W15 agreed that abstract forms of representation, such as “cartoon styled (M5) or “orb-like” voice interface (W14, W15), were perceived as less deceptive and more broadly acceptable. W15 further explained that embodiment affects comfort: “If it’s an avatar, it requires more attention, more effort to be present, like I’m performing extra socially.” Text-based agents, on the contrary, were viewed as lightweight and flexible, more easily integrated into everyday practices such as Slack channels.
@@
Another dimension concerns the level of initiative that Manager Clone Agents should exercise, which participants saw as varying by task. At the lowest level, agents might simply observe or log interactions without intervening. At intermediate levels, they could relay pre-authorized statements or provide clarifications. The highest levels would involve spontaneous contributions or even steering discussions that many participants found unacceptable for sensitive or high-stakes contexts. As noted in subsection 7.2, M16 and M17 worried about being misrepresented by agents making statements they had not endorsed, or by discussions diverted in their absence in ways that later shaped decisions. M18 cautioned that disclosure must be carefully limited, warning, “The authority to disclose information should actually be tightened as well. For example, if it involves company finances and someone says that cash flow is tight, workers might panic.” Similarly, M19 argued that strong presence could backfire sometimes, suggesting, “Don’t make their presence too strong. Even just occasional reminders or emotional support are enough to help everyone gradually accept this new colleague. If their presence is too strong, everyone will feel threatened.”
@@
Finally, the scope of action delegated to Manager Clone Agents emerged as a central concern. Low-delegation configurations would confine agents to communication support, while high-delegation systems could independently execute managerial tasks, such as reallocating resources or authorizing workflows, without human intervention. M13 stressed limits: “As a strictly constrained ‘assistant,’ emphasize ‘reminders’ rather than ‘decisions,’ with transparency and explainability prioritized.” M18 described the risks of overstepping: “Some overly specific suggestions should be constrained. The remaining decision space should be left to the person doing the work. Otherwise, it’s like having a CTO constantly coaching right next to the coder, which makes workers uncomfortable.” M9 proposed explicit labeling: “Label the level of an authorized agent, for example, before getting a Gold level no one pays attention, but after getting Gold any command can be executed. So you can design agent tiers to let the interlocutor understand their current permissions.”
@@
7.4. Toward Disruptive and Unsettling Alternative Futures in Speculative Imagination
While most participant imaginations extended familiar managerial functions, some imagined a set of destabilizing, speculative futures that challenged participants’ own assumptions about hierarchy, authority, and the necessity of human managers altogether.
7.4.1. What if Workers Say “No” to Clone Agents?
Most participants expressed reluctance to use Manager Clone Agents either entirely or in specific situations, yet they also admitted their ability to resist is structurally weak.
Both managers such as M2 and workers such as W1 suggested that open refusal would likely lead to replacement. They agreed on a future where “Those who say ‘no’ will be filtered out and replaced by workers more willing to comply.” W14 added: “It all depends on who has negotiating power. In industries where talent is scarce, workers will successfully refuse.
Despite this pessimism, workers articulated concrete conditions for workplace fairness, including access to training data, a third party supported appeals mechanism, veto rights, etc. W3 emphasized: “The specific reasons behind AI decisions need to be transparent, and we should be informed about what data will be collected about us.” On the other hand, W15 stresses the need for human managers to be involved in some cases: “Decisions involving firing, recruiting, or promoting must be reviewed by a human, as they directly affect people’s livelihoods.” W12 added that a mechanism could be set up to appeal AI decisions, for example, a third party or a labor union could organize collective appeals against decisions. W14 took this speculative future further: “Such veto power would inevitably run into workplace politics that varies between companies. Workers would try to manipulate any veto power they have, and managers would try to use it to exert as much control as they can. Budgets and roles would be questioned, and confusion would come hand in hand with the speed the clone agents bring.
7.4.2. What If Managers Fade?
Several managers including M2, M4, and M5 articulated anxieties that outlined a provocative scenario: a future in which the presence of real managers becomes optional.
This anxiety was captured most vividly in M2’s fiction, which began as a playful thought experiment but drifted, in his words, into “an apocalyptic film.” He sadly imaged, “My clone agent knows their tasks, sometimes better than I do. When I walk by, no one even looks up. They already got what they needed.” Similarly, M4 feared a future where her Manager Clone Agent “gets along with the team better than I do,” absorbing the subtle interpersonal knowledge that she used to define her authority.
The speculative futures raised by managers like M2 and M4 hinted at a more profound philosophical question that cannot be resolved through merely boundary-setting. In M4’s views, if clone agents begin to take over the emotional scaffolding, then the anxiety participants described is not about technological encroachment but a deeper ontological shock: what if leadership identity no longer takes “the unique presence of a person” as a necessary condition, but become an outsource-able “set of functions.” Under this view, M4 believed the challenge is not simply to design against replacement; it is to articulate a new philosophy of management that survives the possibility of replaceability. In this context, it was striking that themes of care, human-centeredness, and relational ethics surfaced repeatedly among managerial participants, as anchors they believed distinguish human leadership from any conceivable clone. As M5 put it, “If the staff only need instructions, then yes, the agent can do that. But if they need to feel understood as people, they still needs me.”
Yet even as participants defended these relational ethics as uniquely human, their speculative futures revealed a more unsettling and liberating possibility: What if organizations themselves did not need managers to flourish? As W7 hinted, a workplace mediated by clone agents might buffer harmful interactions, reduce emotional volatility, and allow workers to focus on meaningful work. W3 frankly admitted: “If the inclusion of AI can diminish, or even overturn power structures like leader–worker dynamics, then as a worker I’m willing to support anything that can facilitate this transformation.” In such visions, the absence of managers is not a loss but a condition for greater happiness, with more room for individuals to exercise agency.
7.4.3. What If Clone Agents Become Tools of Resistance?
Although most workshop discussions centered on how Manager Clone Agents might extend managerial authority, several participants sketched futures in which clone agents are repurposed by workers as instruments of resistance, or even counter-control. These imaginaries surfaced not as fully developed narratives but as hesitant comments and defensive humor.
Workers such as W7 were among the first to hint at this inversion. She desires to use the clone agent as an buffer, shielding herself from emotionally charged interactions. As she put it, “It can help me keep the evidence straight. If the boss says different things on different days, the agent will know.” In many other stories participants envisioned, workers use this upward surveillance to advocate for their own interests. In M10’s story, the team frequently accessing the manager’s agent without notice during the manager’s absence, exposing a series of mistakes that ultimately caused the manager to be demoted. M5 jokingly mirrored this risk: “If my clone keeps a memory of everything I’ve said, it turns into a reflection of me, letting people search for anything they like.” For some workers such as W3, this was not a joke but a glimpse of a more equitable future. “It makes management and evaluation method transparent.” as W3 commented.
A particularly radical scenario emerged in speculative remarks about “turning the system around.” M14 imagined a story where he exploited system transparency. He wrote: “I’ve figured out what Georgebot’s goals are and I can game the system to get my team more funds.” After a year, the worker in his story had move on to another company, because of tech turnover, “I’ll get to say I was trusted with a sizeable budget at a major AI-driven company. Recruiters will see that I worked with AI and they will be impressed. I will continue my embezzlement at my new company, which is also using AI bosses.”
Across these speculative threads, participants believed if Manager Clone Agents become integral to workplace communication and coordination, workers will inevitably develop ways to appropriate, reinterpret, and strategically mobilize these systems.